MONT 106Q – Mathematical Thinking

Study/Discussion Questions on Herodotus, The Histories, Books 1, 2 and 7 September 2016

Note: In the Landmark Herodotus, we will be discussing only

• the Introduction by Rosalind Thomas, (pages ix - xxxvi),

• several sections in Book 1: the Proem (pp. 3), 1.8 - 1.14 (pp. 8 - 10, Gyges and Kandaules), 1.28 - 1.55 (pp. 17 - 30, Solon and Croesus, the death of Croesus’ son Atys, Croesus consults the oracles), 1.84 - 1.92 (pp. 48 - 53, Croesus is defeated but saved on the pyre, his subsequent relations with Cyrus, consideration of the role of the oracles), 1.131 - 1.140 (pp. 71 - 74 religion and customs of the Persians)

• the first section of Book 2 (2.1 - 2.88, pages 117 - 153),

• and most of Book 7 (7.1 - 7.152, pages 493 - 560, then 7.172 - 7.239, pages 569 - 598).

Related to the Introduction:

1. Where and approximately when did Herodotus live? What were the events of the recent past for him that he wanted to preserve a memory of?

2. Why has Herodotus been called the “father of history”? Who originated that name for him? Why has he also been called the “father of lies”? What does Rosalind Thomas say about his reliability as a source of information about the things he describes? What methods did he use to assemble all of his facts and stories?

3. In the Proem, the famous short section at the start of Book 1 describing his purpose in preparing this massive work, Herodotus calls it (in the original Greek) a histori\overline{e}. This word is the source of our word “history,” but Herodotus was essentially writing before what we would call history existed (either as an academic discipline or as a category of written works). What does the Greek word actually mean and what does Rosalind Thomas say about how we should interpret this? How was Herodotus a man of his time and place?

4. How is Herodotus different from the other ancient Greek literature that preceded him (especially the epics of Homer – the Iliad and Odyssey)? How is his writing also different from modern historical works?

5. It is thought that at least sections of the Histories were read aloud by Herodotus at public performances on occasions like the Olympic Games (which included contests related to literature, drama, and music in addition to athletic events in their original Greek format). How might that fact explain some of the choices he made about which topics and stories to include?

Related to our sections in Book 1:

1. What are the events in the story of Gyges and Kandaules? Do you think Kandaules crosses a boundary that should never be crossed there? How is Kandaules’ wife (perhaps surprisingly) in control of the outcome of this episode? What do you think about her actions?

2. Look up Solon (either in previous sections of Herodotus or elsewhere). Who was he and why was he a key figure in Athenian history? What is surprising about the way he interacts with Croesus? What exactly does Solon say about riches and the possibility of living a good life (or having a good death)? Does having a good life mean that you die well?

3. Consultation with oracles plays a huge role in the stories about Gyges, Croesus, and in later portions of Herodotus’ story as well. Look up the Delphic oracle (the priestess who gave the oracle’s pronouncements at Delphi was called the Pythia) – what were the religious and social dimensions of this institution in ancient Greece? How did people use the information an oracle gave them? Was it always clear what the oracle meant? What does Herodotus say about this?

4. What does Herodotus say about the religion and customs of the Persians? Does he seem to have a positive, a negative, or a balanced view of them? How do you suppose other Greeks reacted to his descriptions of the enemies they had fought in a war for national survival? Or was it that? Would the Persians have forced the Greeks to give up their own identity? Is that the way their empire worked?

Related to the first section of Book 2:

Note: I have chosen to have us read only this section of Book 2 because it deals with things Herodotus claims to have observed in person during a visit or visits to Egypt, and it is a good illustration of the sorts of things he found interesting. By way of contrast, the later sections of Book 2 present a version of Egyptian history he claims to have heard from priests in temples he visited. With the knowledge gained by modern archaeology, it can be seen that Herodotus’ accounts are highly inaccurate and misleading. Apparently, at the time Herodotus was writing, no accurate account of the various dynasties, rulers and their deeds had been preserved. This is perhaps not too surprising given the extremely long time span of the ancient Egyptian civilization and its complicated history.

1. What was Psammetichos’ experiment and what did he “discover”? Why do you suppose Herodotus includes this story?

2. What are the possible explanations of the cause of the annual Nile river floods that Herodotus considers and which does he eventually say he supports? (Of course he is totally wrong about this!) What would a modern scientific explanation be? Does this error decrease his believability?

3. What does Herodotus say about the ways Egyptian customs are “completely opposite” those of other peoples?

4. What does he say about Egyptian religious practices? He seems to believe that the gods the Egyptians worship are the same as the Greek gods and he even uses Greek names for them. What evidence does he give for this? Is this believable? Is he being “Hellenocentric” or just naive?

5. What does Herodotus say about the role of cats and crocodiles in Egyptian society? What other specifically Egyptian animals does Herodotus mention?

6. How did the Egyptians treat corpses? What are the embalming and mummification methods they use? (By the way, this section of Herodotus’ account is known to be essentially exactly correct!)

Related to Book 7:

Note: This book presents a central episode in the Greek-Persian wars that occurred in 480 BCE – the invasion of Greece by the Persian king Xerxes and the battle of Thermopylae. This was the encounter between a band of 300 Spartans led by the warrior Leonidas and the much larger Persian invasion force where the Greeks fought a desperate and doomed rearguard action to delay the Persians. The Persians eventually overran this force, and moved on to sack Athens and devastate other areas of Greece. But later that same year the Persian navy was decisively defeated by the Athenians at the battle of Salamis described in Herodotus’ Book 8, and their land army was defeated at the Battle of Plataea the next year. The Persians were forced to withdraw and did not invade Greece again. The Persian empire was eventually completely annihilated by Alexander the Great about 150 years later. This is one of the best-known passages in Herodotus’s story and has formed the basis for a number of portrayals of these events in popular culture, such as in the film 300 from several years ago.

1. In the long chapter 7.8, Xerxes proposes to invade Greece a second time, following on an earlier invasion by his father Darius. Why does he say the Persians should undertake this? He asks two of his lieutenants, Mardonios and Artabanos, to comment and they provide opposite points of view. What are the arguments each of them presents?

2. Does Xerxes seem like a confident and decisive leader here? Why do visions and dreams play such a big role in his decision-making process? How is Artabanos finally won over to the invasion plans? And what’s going on in 7.19? Why does Xerxes seem to overlook such clear negative omen??

3. How long did Xerxes’ invasion preparations take and what steps did he take to ensure success?

4. Some background you need to understand this story – when the Persians sought to conquer a new city or country, they would first send ambassadors to ask whether the people would submit to the rule of the Persian king by offering gifts of earth and water. If the people of the country did so, they were said to medize, after the name of one of the earlier conquests of the Persians, the group called the Medes. Note that many Greek cities actually “medized.” What did Athens and Sparta do when they were asked to “medize”? How was their reaction calculated to infuriate the Persian king? Did he retaliate against the messengers that the Spartans sent to him later? Why or why not?

5. Looking at his great invasion force, in 7.45-46, Xerxes weeps. Why is this?

6. How do the Persians count the number of soldiers they have? What is the tally of all their soldiers Herodotus reports? (Does this seem possible? If you’re interested, Appendix R of the Landmark Herodotus starting on p. 819 attempts to come up with a more realistic estimate of the size of Xerxes’ forces.) Where do they all come from?

Why does Herodotus go into so much detail about their clothing and weapons? (If you have read the Iliad, do you see a parallel with well-known section there?)

7. Who is Demaratos and what is his role in this story?

8. In ancient times, the “judgment of Herodotus” at 7.139 was one of the most controversial sections in the Histories. Why would this infuriate some Greek readers or hearers?

9. Note the repeated consultation of the Delphic oracle in 7.141-142. Who convinced the Athenians that the meaning of the oracle’s second message was that Athens had to rely on its fleet to defeat the Persians? (This is the set-up for the later battle at Salamis mentioned above.)

10. Note that we will skip the sections 7.153 - 7.171 which deal with some side-issues.

11. How does the confrontation at Thermopylae come about? How is it that such a small force of Greeks manages to hold back such a huge Persian army there? What are the physical characteristics of the battle-site that make this possible? How do the Persians eventually overcome the resistance of the Greeks?

12. What are the inscriptions that Herodotus reports on the monuments erected later to the memory of the Greeks who fell at Thermopylae?

13. The one at the bottom of page 593, in particular, has become a sort of exemplar of a warrior’s epitaph emphasizing the nobility of the sacrifices warriors make for their homelands. It has been reused and appropriated in various forms, both with and without irony. Look this up online and find three examples.